40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Wed Sep 3 12:02:19 2014, Tara, Evan, DailyProgress, BEAT, Beat found 
    Reply  Wed Sep 3 16:34:37 2014, Tara, Evan, DailyProgress, BEAT, Beat, mode-matching noiseBudget.pdf
Message ID: 1500     Entry time: Wed Sep 3 16:34:37 2014     In reply to: 1499
Author: Tara, Evan 
Type: DailyProgress 
Category: BEAT 
Subject: Beat, mode-matching 

Quote:

Tara added some more juice to the north cavity heater last night. Now we can lock both cavities to TEM00 and get a beat within the bandwidth of the 1811.

  • North laser slow: 5.020 V
  • South laser slow: 0.722 V
  • Beat frequency: 49.3 MHz

Beat frequency drifted to 61 MHz over the course of a few hours. We need to wait for the cavity temperatures to settle.

I improved the mode-matching a little bit on the south cavity; it's about 50% (the theoretical max is 71%). The south lenses are now on translation stages.

I've attached a beat spectrum. Nothing is floated, RAM is not optimized, etc.; this is just a rough indicator of where things stand.

Here is what I think should happen next, in rough order of importance:

  1. Float chamber
  2. Measure RIN
  3. Measure photothermal TF (I also need to recheck my photothermal code — I don't believe the coating TE part)
  4. Put photothermal noise on noise budget
  5. Reduce RAM.
  6. Measure residual frequency noise.
  7. Measure PLL noise. Use ATF DAQ and make spectral histogram.
  8. Measure seismic noise (with Guralp or T240), with table floated and unfloated. Use ATF DAQ and make spectral histogram.
Attachment 1: noiseBudget.pdf  1001 kB  | Hide | Hide all
noiseBudget.pdf
ELOG V3.1.3-